# On Large Language Models in Mission-Critical IT Governance: Are We Ready Yet?

### v1

@@ IV Discussions
Implications for Researchers. Our findings suggest that researchers have a crucial role in addressing the limitations and ethical challenges associated with LLM deployment in mission-critical contexts. One primary area for further research is the development of context-sensitive LLMs, i.e., regulation-oriented models (Q16). Our previous attempts showed that RAG improved contextual knowledge, reducing hallucinations in RA [10], yet future efforts will be focused on investigating means to realize more regulations-oriented LLMs. For instance, due to the current state of the art, no ISO standards or national and international laws are not algorithmic-friendly regarding how an algorithm, or AI approach, can tackle and operate with zero to no deviation from them. Similarly, LLM-based governance is a branch of research that is not currently receiving substantial research effort. Our work highlighted many aspects, from regulation compliance (Q16) to missing ethical framework, the needs for accountable, explainable, and transparent AIs and its reliability (Q17).
@@
Open-ended answer to Q16 also highlights that the practitioner’s unexpected confidence towards LLMs is deemed possible only as long as we do not “rely exclusively on it” meaning that LLMs need to be mere tools in the hand of human experts, not replacing them. Practitioners evidenced that LLMs “might not take into account some variables that are fundamental in the analysis of various problems”. Transparency and accountability are also key concerns; users should prefer systems that clearly explain how models arrive at their decisions and allow auditing to ensure the systems meet accountability standards (Q30). Those concerns are highlighted in one answer to the open part of Q17 that state: “LLMs are often considered ‘black boxes,’ making it difficult to understand how they arrive at their conclusions.”, thus highlighting the daunting need of explainable AIs. In the same vein, a specific practitioner answer to the open-ended part of Q22The introduction of an LLM in a company has the potential to have a great positive impact and it is possible to expect excellent results, but its continuous updating and training is essential to continue to improve it, and not all companies can sustain such an organizational/economic impact” summaries the positive, yet caution approach that is evident when analyzing our collected data. Finally, a major impact emerge as the possibility of having a tool that harmonizes technical requirements, i.e. driven solely by the internal logic of the industry, with regulatory needs. From an industrial perspective, practitioners highlighted the need for a shared tool, perhaps developed by organizations like ACM/IEEE, to benchmark software compliance along with regulatory experts, both national and international, in the development process of such tools.
@@
Implications for Policy Makers Our findings show that policymakers should focus on defining a clear framework for LLMs used in Mission-Critical IT Systems, establishing the rules to keep them current, and monitoring their use, as deduced from Table VII. The ethical landscape is crucial for practitioners. Supporting regular ethics training and promoting collaboration in policy-making between researchers, policymakers, and industry stakeholders will help foster an environment where all the specific knowledge of each actor involved can be leveraged to craft a state=of-the-art legislation that will guide an ethical and safe integration of GAIs in GMCS (Q25, Q31). Practitioners advocate for standards and regulations that mandate regular audits (Q27), encourage interdisciplinary cooperation, and align with established safety and accountability practices. By actively engaging in these initiatives, practitioners contribute to a future where LLMs can sustainably and ethically augment mission-critical governance processes. Global AI regulation is fragmented. The EU has its AI Act, while U.S. legislation is still developing. China’s AI law is in draft form [37], and Russia has an AI Framework running from 2020 to 2025 [38]. The UN adopted a 2024 resolution promoting “safe, secure, and trustworthy” AI, but without a clear global consensus [39]. There is no unified vision or international body to oversee AI development, ethics, and security, leading to inconsistent governance across regions.
Focusing on the perimeter of the EU, it is paramount to accelerate the definition of a more technical framework (Q28) by hosting regular workshops and seminars with stakeholders (Q32), establishing both a permanent public-private partnership for collaborative research (Q32) and an advisory board (Q32), with representatives from various sectors, to respond to any technical (Q30 and Q31), security and ethical concerns (Q31 and Q33) or questions related to privacy (Q26 and Q30) that could arise. Specifically, an answer to the open part of Q30 states that “The use of these tools should result from an accreditation process similar to that required for handling classified security information”. Such committees should define strict technical rules and guidelines (Q31), define precise documentation requirements for LLM algorithm and training data (Q27) as well as develop industry-specific guidelines for LLM usage and establish a certification process for LLM technologies (Q27) and have the powers to enforce regular audits and compliance checks for LLM deployments (Q27). Finally, the creation of interdisciplinary committees to draft comprehensive policies benchmark of policies (Q33) and the benchmarking of policies from other countries that use LLMs will also be fundamental to keep them up to date and to address them with better solutions adopted elsewhere (Q33).
@@ V Threats to Validity
External Validity concerns how the research elements (subjects, artifacts) represent actual elements [14]. Our sample size, though not as large as another similar study, is justified by the niche context the company operates in and is still smaller than those reflecting all possible contexts of IT MCS practitioners. We tried to mitigate this threat by increasing the diversity of the participants in organizational size, expertise, and experience and geographical location. Nonetheless, due to the referenced process and regulations being primarily international, our findings may represent other countries, though considering the implementation allowance the international laws grant to each nation. Given these limitations, we plan to survey with overseas partners to synthesize results across time and the development of a more robust, empirically supported understanding.
Another potential threat is the lack of control over the sample selection, which may cause the sample to be biased toward individuals who are very interested or biased toward the subject under study. This may, hence, outweigh the positive or negative perception results. Nevertheless, we found that a considerable proportion of participants at baseline mainly had cautious positive perceptions, so any negative bias is probably slight.
@@ VI Related Works
Scientific works regarding LLMs and their possible impact on research fields and industrial use cases have started rising prominently on our field landscape [10, 3, 40, 41, 42]. LLMs currently impact all scientific fields, which can leverage a ’copilot.’ From healthcare to software development and security, LLMs are among the pinnacles of our research achievements [40, 3].
For instance, Fan et al. [40], surveyed the emerging area of LLMs in Software Engineering (SE) and identified open research challenges related to their application. They highlighted the novelty and creativity LLMs bring to various SE activities but also pointed out the technical challenges, such as the need to filter out incorrect solutions. The survey emphasized the crucial role of hybrid techniques combining traditional SE methods with LLMs to develop reliable and effective LLM-based SE solutions.
Futhermore, Kirova et al. [41], discussed the growing significance of generative AI, particularly LLMs, in various sectors and their impact on software engineering. They emphasized the need for software engineering education to adapt to the emerging LLM environment, addressing a gap in the literature regarding LLMs’ specific implications for this field. The paper explored the goals of software engineering education, changes to the discipline, course pedagogy, and ethics, arguing for a holistic approach that combines technical skills, ethical awareness, and adaptable learning strategies.
In the same vein as the general cautious approach towards these new advances, Sallou et al. [42] highlighted the significant impact of LLMs on various SE tasks, such as code completion, test generation, program repair, and code summarization. Despite their potential, the authors cautioned researchers about the complex factors affecting the outcomes of LLM-based experiments. The paper discussed potential threats to LLM research’s validity, including issues like closed-source models, data leakage, and reproducibility challenges. In response, the authors proposed guidelines for SE researchers and LM providers to address these concerns, demonstrating the guidelines’ implications through existing best practices and a practical example in test case generation.
Differently from the broader survey and analysis provided by fellow researchers, the context of our study is, indeed, quite peculiar. The usual perception of new or trending technologies is negative or, worst case, desk-rejected. Recently, Esposito et al. [10, 3] has tackled a primer for LLMs in MC risk analysis, building a cornerstone on which fellow researchers and industrial practitioners can build.
Inspired by their preliminary results, we designed the present survey to grasp practitioners’ perceived benefits and fears in the adoption at large in IT MCS governance.

### v2

@@ IV Discussions
Implications for Researchers. Our findings suggest that researchers have a crucial role in addressing the limitations and ethical challenges associated with LLM deployment in mission-critical contexts. For instance, in a recent work, Shi et al. [40] explored LLMs’ distractibility, i.e., the tendency for a model to be distracted by irrelevant context. Human experts analysing LLMs output in our past study, highlighted that a small detail in the description of a scenario led the model to hallucinate and make up non-existing threat tied to the small detail [10] thus confirming Shi et al. work. Therefore, a primary area for further research is the development of context-sensitive LLMs, i.e., regulation-oriented models (Q16).
Our previous attempts showed that RAG improved contextual knowledge, reducing hallucinations in RA [10], yet future efforts will be focused on investigating means to realize more regulations-oriented LLMs. For instance, due to the current state of the art, no ISO standards or national and international laws are not algorithmic-friendly regarding how an algorithm, or AI approach, can tackle and operate with zero to no deviation from them. Similarly, LLM-based governance is a branch of research that is not currently receiving substantial research effort. Our work highlighted many aspects, from regulation compliance (Q16) to missing ethical framework, the needs for accountable, explainable, and transparent AIs and its reliability (Q17).
@@
Open-ended answer to Q16 also highlights that the practitioner’s unexpected confidence towards LLMs is deemed possible only as long as we do not “rely exclusively on it” meaning that LLMs need to be mere tools in the hand of human experts, not replacing them. Practitioners evidenced that LLMs “might not take into account some variables that are fundamental in the analysis of various problems”. Transparency and accountability are also key concerns; users should prefer systems that clearly explain how models arrive at their decisions and allow auditing to ensure the systems meet accountability standards (Q30). Those concerns are highlighted in one answer to the open part of Q17 that state: “LLMs are often considered ‘black boxes,’ making it difficult to understand how they arrive at their conclusions.”, thus highlighting the daunting need of explainable AIs. In the same vein, a specific practitioner answer to the open-ended part of Q22The introduction of an LLM in a company has the potential to have a great positive impact and it is possible to expect excellent results, but its continuous updating and training is essential to continue to improve it, and not all companies can sustain such an organizational/economic impact” summarizes the positive, yet caution approach that is evident when analyzing our collected data. Finally, a major impact emerge as the possibility of having a tool that harmonizes technical requirements, i.e. driven solely by the internal logic of the industry, with regulatory needs. From an industrial perspective, practitioners highlighted the need for a shared tool, perhaps developed by organizations like ACM/IEEE, to benchmark software compliance along with regulatory experts, both national and international, in the development process of such tools.
@@
Implications for Policy Makers Our findings show that policymakers should focus on defining a clear framework for LLMs used in Mission-Critical IT Systems, establishing the rules to keep them current, and monitoring their use, as deduced from Table VII. The ethical landscape is crucial for practitioners. Supporting regular ethics training and promoting collaboration in policy-making between researchers, policymakers, and industry stakeholders will help foster an environment where all the specific knowledge of each actor involved can be leveraged to craft a state-of-the-art legislation that will guide an ethical and safe integration of GAIs in GMCS (Q25, Q31). Practitioners advocate for standards and regulations that mandate regular audits (Q27), encourage interdisciplinary cooperation, and align with established safety and accountability practices. By actively engaging in these initiatives, practitioners contribute to a future where LLMs can sustainably and ethically augment mission-critical governance processes. Global AI regulation is fragmented. The EU has its AI Act, while U.S. legislation is still developing. China’s AI law is in draft form [41], and Russia has an AI Framework running from 2020 to 2025 [42]. The UN adopted a 2024 resolution promoting “safe, secure, and trustworthy” AI, but without a clear global consensus [43]. There is no unified vision or international body to oversee AI development, ethics, and security, leading to inconsistent governance across regions.
Focusing on the perimeter of the EU, it is paramount to accelerate the definition of a more technical framework (Q28) by hosting regular workshops and seminars with stakeholders (Q32), establishing both a permanent public-private partnership for collaborative research (Q32) and an advisory board (Q32), with representatives from various sectors, to respond to any technical (Q30 and Q31), security and ethical concerns (Q31 and Q33) or questions related to privacy (Q26 and Q30) that could arise.
Experts in our survey and pilot study expressed concerns about the limited reasoning capabilities of LLMs. This perspective aligns with findings by Mirzadeh et al.[44], who evaluated LLMs’ mathematical reasoning using GSM8K and introduced GSM-Symbolic for more transparent and explainable assessments. Their study revealed significant performance drops when numerical values or question phrasing changed, highlighting the models’ reliance on pattern replication rather than genuine reasoning. These findings, combined with expert feedback, drive our motivation to explore Large Reasoning Models (LRMs), such as OpenAI’s o1 and o1 preview, or open-source alternatives like Marco-o1[45], i.e., model optimized for outputting alongside their response their inner reason, to address this limitation. In the same vein, when clarity and transparency is threatened, experts suggested in the open part of Q30 that “The use of these tools should result from an accreditation process similar to that required for handling classified security information”. Future research and industrial partnerships should establish a foundation for an accreditation process that ensures LLMs transparently disclose their inner reasoning, enabling thorough evaluation and validation as well defining strict technical rules and guidelines (Q31), define precise documentation requirements for LLM algorithm and training data (Q27). Moreover such a collaboration should also develop industry-specific guidelines for LLM usage and establish a certification process for LLM technologies (Q27) and have the powers to enforce regular audits and compliance checks for LLM deployments (Q27). Finally, the creation of interdisciplinary committees to draft comprehensive policies benchmark of policies (Q33) and the benchmarking of policies from other countries that use LLMs is also paramount to keep them up to date and to address them with better solutions adopted elsewhere (Q33).
@@ V Threats to Validity
External Validity concerns how the research elements (subjects, artifacts) represent actual elements [14]. The sample size is relatively small compared to studies on the same topic but is justified by the niche context of GMCS practitioners. To address this limitation we increased the diversity of participants by varying their organizational size, expertise, experience, and geographical location. Since the laws and regulations considered in our study is mostly international, our findings is relevant for other countries within the bounds of such laws and standards. Future works will focus on collaborating with overseas partners to carry out empirical studies leading to a deeper understanding of the potential and limitations of LLMS in GMCS. Moreover lack of control in selecting the samples constitute a threat, which might introduce unbalanced bias from participants with strong opinions. Nevertheless, we found that a considerable proportion of participants were sitting on the fence, with a cautious positive perceptions, so the probability of a “strong opinions” bias is probably slight.
@@ VI Related Works
The scientific landscape is witnessing a surge in research on the impact of LLMs across various domains, from healthcare to software development and security [10, 3, 46, 47, 48]. These models, treated as “copilots of the innovation” [10], represent a pinnacle of research innovation, offering significant potential to advance both academic and industrial applications [46, 3].
For instance, Fan et al. [46], explored the emerging use of LLMs in Software Engineering, identifying both their potential for enhancing creativity and novelty in SE tasks and the technical challenges, such as filtering incorrect solutions. They emphasized the importance of hybrid approaches that integrate traditional SE methods with LLMs to develop reliable and effective solutions.
In the same vein as the general cautious approach towards these new advances, Sallou et al. [48] explored LLM applications in SE tasks and raised concerns about validity threats, including data leakage and reproducibility. They proposed guidelines for researchers and LLM providers, illustrating their use with a test case generation example.
Differently from past studies we focused on a peculiar context. We recently tackled a primer for LLMs in MCs RA [10, 3], building a cornerstone on which fellow researchers and industrial practitioners can build. To deepen our investigation, we designed the present survey to grasp practitioners’ perceived benefits and fears in the adoption at large in IT MCS governance.
@@ VII Conclusions
Acknowledgment
This work has been partially funded by projects Business Finland 6GSoft, and Research Council of Finland MuFAno (grants n. 349487 and 349488).